OpenStudy (anonymous):

Need some help please.. Obtain the general solution: y'(y''-y)=x Need help on the Yp solution..

OpenStudy (anonymous):

The problem is shown as: \[(D^3-D)y=x\] Its undetermined coefficients and I need help finding the Yp solution..

OpenStudy (anonymous):

Would Yp=A+Bx ?

OpenStudy (accessdenied):

this is confusing me: (D^3 - D)y = x Is this notation such that: \( Dy = \dfrac{d}{dx} y \) \(D^3 y = \dfrac{d^3}{dx^3} y \) idk why this is throwing me off so much lol

OpenStudy (anonymous):

Lol i believe it's an old form of notation.. I think it means the same as y'(y''-y)=x

OpenStudy (accessdenied):

I believe you distribute it through, like this: \( (D^3 - D) y = D^3 y - Dy = \dfrac{d^3}{dx^3} y - \dfrac{dy}{dx} \)

OpenStudy (accessdenied):

if you factored it, you'd have to "create" a y to go with both factors really.

OpenStudy (anonymous):

Ok, gotcha.. now im getting my Yc as: \[Yc=C1+C2e^{x}+C3e^{-x}\]

OpenStudy (accessdenied):

hmm, where'd you get that from?

OpenStudy (anonymous):

For the left hand side only.. \[(D^3-D)y=0\] \[e^{mx}(m^3-m)=0\] \[m(m^2-1)=0\] \[m(m-1)(m+1)=0\] \[m=0,1,-1\] \[Solution: e^{0x},e^x,e^{-x}\] \[Yc=C1+C2e^x+C3e^{-x}\]

OpenStudy (accessdenied):

ohh you got it, was thinking m^3 - 1 but its Dy not y. :p so Y_c looks good

OpenStudy (accessdenied):

for Y_p, I'm thinking about how the equation plays out: third derivative of polynomail - first derivative of polynomial = x The smallest order derivative is that first order, and so we should test something whose first derivative is linear i am thinking.

OpenStudy (accessdenied):

the third degree isn't quite important because our only interest is matching the right-hand side's linear term, x. as long as our first derivative contains an x term, we should be fine at getting a match. anything residual will be 0.

OpenStudy (anonymous):

I see... thats why im thinking that Yp would = A+Bx, but it doesnt look right when i try to plug it in...

OpenStudy (accessdenied):

well, Y_p = A + Bx contains a linear term, but its first derivative is a constant, right?

OpenStudy (anonymous):

Exactly... So that would lead me to think that: \[Yp=Ax^2+Bx+C\]

OpenStudy (accessdenied):

i think that will work out better for you. :)

OpenStudy (anonymous):

\[Yp=\frac{ 1 }{ 2 }x^2\]?

OpenStudy (accessdenied):

i think you have one small sign to recheck there

OpenStudy (accessdenied):

(third derivative = 0) minus (first derivative = 2A x + B) = x

OpenStudy (anonymous):

You guys are on the right path, are you familiar with reduction of order, ie for example just let v(x) = y'(x)? (I believe guys are basically doing the same thing, but don't take my word for it)

OpenStudy (accessdenied):

oh, i see what you're getting at @Crixpack from y' ' ' - y ' = x <-- just reducing the order here with v = y' to v ' ' - v = x and then finding v there / referring back to v = y' to find y?

OpenStudy (anonymous):

Your gold @AccessDenied !!! Got it!

OpenStudy (anonymous):

Yup, exactly, and you get the same answer you guys do. Cheers :)

OpenStudy (accessdenied):

good work :) and i do believe you could reduce order as well, in fact i hadn't even noticed it until it was pointed out! but either way i think you get the same answer. :)